OpenStudy (anonymous):

What is the sum of a 28-term arithmetic sequence where the first term is 12 and the last term is 255?

OpenStudy (anonymous):

Formula: Sn = n(a1 +an)/2 So, \[S28= 28 ( 12 +255)/2\] Simplify

OpenStudy (anonymous):

3318

OpenStudy (anonymous):

2,937 3,204 3,471 3,738

OpenStudy (anonymous):

@TwinYang

OpenStudy (anonymous):

You didn't simplify correctly, simplify again.

OpenStudy (anonymous):

got the same thing

OpenStudy (anonymous):

Simplify 12+255 to 267 x=28×267/2 2 Simplify 28×267 to 7476 x=7476/2 3 Simplify 7476/2 to 3738 x=3738

OpenStudy (anonymous):

oh thank you!